To create a breakfast beverage, pineapple juice in two concentrations, 16% and 30%, must be combined into a solution that will be mixed with another type of juice to produce the beverage. If 50 gallons of 16% juice is used, how many gallons of the 30% juice must be used to obtain a 20% pineapple juice solution?
Answer by mananth(16946)

To create a breakfast beverage, pineapple juice in two concentrations, 16% and
30%, must be combined into a solution that will be mixed with another type of juice to produce the beverage. If 50 gallons of 16% juice is used, how many gallons of the 30% juice must be used to obtain a 20% pineapple juice solution?
Juice A 16% quantity 50 gal
Juice B 30 % quantity x gal
Total quantity = x+50 gal 20%
Concentration of Juice A in the mixture = 0.16*50
Concentration of Juice B in the mixture = 0.3 x
Concentration of mixture = 0.2(x+50)
0.16*50+0.3x= 0.2(x+50)
8 +0.3x = 0.2x + 10
0.1x= 2
x= 2/0.1
x= 20 gallons Juice B required